### v1

@@ 5 Algorithms and Proofs
Let $\ell^{\textrm{precision}}(g,x)$ and $\ell^{\textrm{recall}}(g,x)$ denote the precision loss and recall loss of graph $g$ at node $x$:
$\ell^{\textrm{precision}}(g,x)=\frac{|N_{g}(x)\setminus N_{g^{\text{target}}}(x)|}{n_{g}(x)}\,,$
$\ell^{\textrm{recall}}(g,x)=\frac{|N_{g^{\text{target}}}(x)\setminus N_{g}(x)|}{n_{g^{\text{target}}}(x)}\,.$
Let ^$\widehat{\ell}^{\textrm{precision}}(g)=\frac{1}{m}\sum\ell^{\textrm{precision}}(g,x_{i})$ and ^$\widehat{\ell}^{\textrm{recall}}(g)=\frac{1}{m}\sum\ell^{\textrm{recall}}(g,x_{i})$ denote the empirical precision and recall losses. Let $a\wedge b:=\min(a,b)$.
@@ 5.1 Maximum Likelihood Method in the Realizable Case
In the realizable setting, the target graph $g^{\text{target}}$ is in the hypothesis class. Given the IID training data $(x_{1},v_{1}),\ldots,(x_{m},v_{m})$, the maximum likelihood method returns the graph
$g^{\text{output}}=\operatorname*{arg\,max}_{g\in\mathcal{H}}\prod_{i=1}^{m}\frac{\mathds{1}(v_{i}\in N_{g}(x_{i}))}{n_{g}(x_{i})}\,.$
In other words, $g^{\text{output}}$ is a graph in $\mathcal{H}$ satisfying
  • consistency: $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin N_{g^{\text{output}}}(x_{i}))=0$
  • regulation: among all consistent graphs, $g^{\text{output}}=\operatorname*{arg\,min}_{g:g\text{ is consistent}}\sum_{i=1}^{m}\log(n_{g}(x_{i}))$.555the base of $\log$ in this work is $2$.
@@
Proof of Theorem 3  It suffices to prove that for any fixed $(x_{1},\ldots,x_{m})$, when $m\geq\frac{12\log(4|\mathcal{H}|/\delta)}{\varepsilon}$, w.p. at least $21-\delta/$ 2 over $v_{1:m}$, the empirical values of recall and precision are small, ^$\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon/2$ and ^$\widehat{\ell}^{\textrm{precision}}(g^{\text{output}})\leq\varepsilon/2$. Then we can show $\ell^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon$ and $\ell^{\textrm{precision}}(g^{\text{output}})\leq\varepsilon$ by applying empirical Bernstein bounds to both precision and recall losses. Now we prove this statement.
Bounding recall loss is easy as $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin N_{g}(x_{i}))$ is an unbiased estimate of recall loss. With probability $41-\delta/$ 4 over the randomness of $v_{i}\sim\mathrm{Unif}(N_{g^{\text{target}}}(x_{i}))$ for all $i\in[m]$, the empirical loss for recall is
^$\widehat{\ell}^{\textrm{recall}}(g)=\frac{1}{m}\sum_{i=1}^{m}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}\leq\sum_{i=1}^{m}\mathds{1}(v_{i}\notin N_{g}(x_{i}))+\varepsilon/6=\varepsilon/6\,,$ (3)
for all consistent $g$ with $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin N_{g}(x_{i}))=0$. Hence, w.p. $41-\delta/$ 4, ^$\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon/6$.
@@
^$\widehat{\ell}^{\textrm{precision}}(g)$
$=$ $\frac{1}{m}\sum_{i=1}^{m}\frac{|N_{g}(x_{i})\setminus N_{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})}$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\frac{|N_{g}(x_{i})\setminus N_{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})}+\frac{1}{m}\sum_{i=1}^{m}\mathds{1}(i\notin A_{g})$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\min\left(\frac{2|N_{g}(x_{i})\setminus N_{g^{\text{target}}}(x_{i})|}{n_{g^{\text{target}}}(x_{i})},1\right)+\frac{2}{m}\sum_{i\notin A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$=$ $\frac{2}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})+|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|-n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})},\frac{1}{2}\right)+\frac{2}{m}\sum_{i\notin A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\frac{2}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}-1,\frac{1}{2}\right)+\frac{2}{m}\sum_{i\in A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}+\frac{2}{m}\sum_{i\notin A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
@@
when $g$ satisfies $\sum_{i=1}^{m}\log(n_{g}(x_{i}))\leq\sum_{i=1}^{m}\log(n_{g^{\text{target}}}(x_{i}))$. Hence, for any graph $g$ with ^$\widehat{\ell}^{\textrm{precision}}(g)>\frac{\varepsilon}{2}$ and ^$\widehat{\ell}^{\textrm{recall}}(g)\leq\frac{\varepsilon}{6}$, we have
@@
The probability of outputting such a graph $g$ is
@@
Hence, with probability at least $21-\delta/$ 2 over $v_{1:m}$, $g^{\text{output}}$ will satisfy
^$\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon/6\,,\widehat{\ell}^{\textrm{precision}}(g^{\text{output}})\leq\varepsilon/2\,.$
@@ 5.2 Modified Maximum Likelihood Method in the Agnostic Case
In the agnostic setting, we shift our goal from finding a graph with nearly zero precision and recall losses to determining, given any $p,r\in[0,1]$, the minimum precision loss $p^{\prime}$ such that $(p,r)\Rightarrow(p^{\prime},r)$.
In fact, we can show something stronger: we do not need to know $p$. Specifically, given any $r\in[0,1]$, let $p=\min_{g:\ell^{\textrm{recall}}(g)\leq r}\ell^{\textrm{precision}}(g)$ be the optimal precision loss among all graphs with recall loss at most $r$. What is the smallest $p^{\prime}$ such that we achieve $\ell^{\textrm{recall}}(g^{\text{output}})\leq r$ and $\ell^{\textrm{precision}}(g^{\text{output}})\leq p^{\prime}$?
@@
Recall that in the realizable case, the maximum likelihood method selects the consistent graph with the smallest empirical log degree, i.e., $g^{\text{output}}=\operatorname*{arg\,min}_{g:g\text{ is consistent}}\sum_{i=1}^{m}\log(n_{g}(x_{i}))$. Basically, the consistency guarantees small recall loss and the empirical log degree is used as a regulation term to bound precision loss. In the agnostic setting, the maximum likelihood method fails as there may be no consistent graph. We present a modified version of the maximum likelihood method, which still has the “consistency” component and adopts log degree as a regulation term. The algorithm operates as follows.
@@
    Finding a set of plausible graphs which make at most $m\cdot(r+\varepsilon)$ mistakes: For any graph $g$, let $I_{g}=\{i|v_{i}\notin N_{g}(x_{i})\}$ denote the indices of training points that the graph $g$ is inconsistent with. Then let ^=$\widehat{\mathcal{H}}=\{g\in\mathcal{H}||I_{g}|\leq m\cdot(r+\varepsilon)\}$ be the set of graphs making at most $m\cdot(r+\varepsilon)$ mistakes.
@@
    Returning the graph with low empirical log truncated degree: Return
@@
Given any $r\in[0,1]$ and $m\geq O(\frac{\log(|\mathcal{H}|)+\log(1/\delta)}{\varepsilon^{2}})$ IID training data, the modified maximum likelihood method can return a graph satisfying
@@
Proof Sketch  Let $g^{\dagger}$ to be the graph with recall loss at most $r$ and precision loss $p$. Let $\overline{r}=r+2\varepsilon$ and $\overline{p}=p+2\varepsilon$. Similar to the realizable setting, we build connection between precision and recall with empirical log degree (Lemma 2), i.e., for any constant $c\in(0,1]$,
^$\widehat{\ell}^{\textrm{precision}}(g)\leq\frac{1+c}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+\frac{1+c}{c}\widehat{\ell}^{\textrm{recall}}(g)\,.$
@@
where $B=\{i|\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}\}$. The term (b) is lower bounded by the recall loss in Lemma 4. Intuitively, if $\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is small, the recall loss must be large while any graph in $\widehat{\mathcal{H}}$ has a small empirical recall loss.
For term (a), since ^$\widehat{\ell}^{\textrm{recall}}(g^{\dagger})\leq\overline{r}$ and ^$\widehat{\ell}^{\textrm{precision}}(g^{\dagger})\leq\overline{p}$, at most training points, $\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is in $[\frac{1}{2},2]$ (if it’s too large at $x_{i}$, precision loss is large; if it’s too small, recall loss is large). Also, for any graph in $\widehat{\mathcal{H}}$, the empirical recall loss is small and thus, at most training points, $\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}$. At these training points satisfying $\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is in $[\frac{1}{2},2]$ and $\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}$, we have
$\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}=\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}+\log\frac{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$

### v2

@@ 5 Algorithms and Proofs
Let $\ell^{\textrm{precision}}(g,x)$ and $\ell^{\textrm{recall}}(g,x)$ denote the precision loss and recall loss of hypothesis $g$ at input $x$:
$\ell^{\textrm{precision}}(g,x)=\frac{|g(x)\setminus g^{\text{target}}(x)|}{n_{g}(x)}\,$,
$\ell^{\textrm{recall}}(g,x)=\frac{|g^{\text{target}}(x)\setminus g(x)|}{n_{g^{\text{target}}}(x)}\,.$
Then the empirical precision and recall losses are $\widehat{\ell}^{\textrm{precision}}(g)=\frac{1}{m}\sum\ell^{\textrm{precision}}(g,x_{i})$ and $\widehat{\ell}^{\textrm{recall}}(g)=\frac{1}{m}\sum\ell^{\textrm{recall}}(g,x_{i})$. Let $a\wedge b:=\min(a,b)$.
@@ 5.1 Maximum Likelihood Method in the Realizable Case
In the realizable setting, the target hypothesis $g^{\text{target}}$ is in the hypothesis class. Given the IID training data $(x_{1},v_{1}),\ldots,(x_{m},v_{m})$, the maximum likelihood method returns the hypothesis
$g^{\text{output}}=\operatorname*{arg\,max}_{g\in\mathcal{H}}\prod_{i=1}^{m}\frac{\mathds{1}(v_{i}\in g(x_{i}))}{n_{g}(x_{i})}\,.$
In other words, $g^{\text{output}}$ is a hypothesis in $\mathcal{H}$ satisfying
  • consistency: $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin g^{\text{output}}(x_{i}))=0$;
  • regulation: among all consistent hypotheses, the output hypothesis satisfies that $g^{\text{output}}=\operatorname*{arg\,min}_{g:g\text{ is consistent}}\sum_{i=1}^{m}\log(n_{g}(x_{i}))$.555the base of $\log$ in this work is $2$.
@@
Proof of Theorem 3 It suffices to prove that for any fixed $(x_{1},\ldots,x_{m})$, when $m\geq\frac{12\log(4|\mathcal{H}|/\delta)}{\varepsilon}$, w.p. at least $1-\delta/2$ over $v_{1:m}$, the empirical values of recall and precision are small, $\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon/2$ and $\widehat{\ell}^{\textrm{precision}}(g^{\text{output}})\leq\varepsilon/2$. Then we can show $\ell^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon$ and $\ell^{\textrm{precision}}(g^{\text{output}})\leq\varepsilon$ by applying empirical Bernstein bounds to both precision and recall losses. Now we prove this statement.
Bounding recall loss is easy as $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin g(x_{i}))$ is an unbiased estimate of recall loss. With probability $1-\delta/4$ over the randomness of $v_{i}\sim\mathrm{Unif}(g^{\text{target}}(x_{i}))$ for all $i\in[m]$, the empirical loss for recall is
$\widehat{\ell}^{\textrm{recall}}(g)=\frac{1}{m}\sum_{i=1}^{m}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}\leq\sum_{i=1}^{m}\mathds{1}(v_{i}\notin g(x_{i}))+\varepsilon/6=\varepsilon/6\,$, (3)
for all consistent $g$ with $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin g(x_{i}))=0$. Hence, w.p. $1-\delta/4$, $\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon/6$.
@@
$\widehat{\ell}^{\textrm{precision}}(g)$
$=$ $\frac{1}{m}\sum_{i=1}^{m}\frac{|g(x_{i})\setminus g^{\text{target}}(x_{i})|}{n_{g}(x_{i})}$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\frac{|g(x_{i})\setminus g^{\text{target}}(x_{i})|}{n_{g}(x_{i})}+\frac{1}{m}\sum_{i=1}^{m}\mathds{1}(i\notin A_{g})$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\min\left(\frac{2|g(x_{i})\setminus g^{\text{target}}(x_{i})|}{n_{g^{\text{target}}}(x_{i})},1\right)+\frac{2}{m}\sum_{i\notin A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$=$ $\frac{2}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})+|g^{\text{target}}(x_{i})\setminus g(x_{i})|-g^{\text{target}}(x_{i})}{n_{g^{\text{target}}}(x_{i})},\frac{1}{2}\right)+\frac{2}{m}\sum_{i\notin A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\frac{2}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}-1,\frac{1}{2}\right)+\frac{2}{m}\sum_{i\in A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}+\frac{2}{m}\sum_{i\notin A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
@@
when $g$ satisfies $\sum_{i=1}^{m}\log(n_{g}(x_{i}))\leq\sum_{i=1}^{m}\log({n_{g^{\text{target}}}(x_{i})})$. Hence, for any hypothesis $g$ with $\widehat{\ell}^{\textrm{precision}}(g)>\frac{\varepsilon}{2}$ and $\widehat{\ell}^{\textrm{recall}}(g)\leq\frac{\varepsilon}{6}$, we have
@@
The probability of outputting such a hypothesis $g$ is
@@
Hence, with probability at least $1-\delta/2$ over $v_{1:m}$, $g^{\text{output}}$ will satisfy
$\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\leq\varepsilon/6\,,\widehat{\ell}^{\textrm{precision}}(g^{\text{output}})\leq\varepsilon/2\,.$
@@ 5.2 Modified Maximum Likelihood Method in the Agnostic Case
In the agnostic setting, we shift our goal from finding a hypothesis with nearly zero precision and recall losses to determining, given any $p,r\in[0,1]$, the minimum precision loss $p^{\prime}$ such that $(p,r)\Rightarrow(p^{\prime},r)$.
In fact, we can show something stronger: we do not need to know $p$. Specifically, given any $r\in[0,1]$, let $p=\min_{g:\ell^{\textrm{recall}}(g)\leq r}\ell^{\textrm{precision}}(g)$ be the optimal precision loss among all hypotheses with recall loss at most $r$. What is the smallest $p^{\prime}$ such that we achieve $\ell^{\textrm{recall}}(g^{\text{output}})\leq r$ and $\ell^{\textrm{precision}}(g^{\text{output}})\leq p^{\prime}$?
@@
Recall that in the realizable case, the maximum likelihood method selects the consistent hypothesis with the smallest empirical log output size, i.e., $g^{\text{output}}=\operatorname*{arg\,min}_{g:g\text{ is consistent}}\sum_{i=1}^{m}\log(n_{g}(x_{i}))$. Basically, the consistency guarantees small recall loss and the empirical log output size is used as a regulation term to bound precision loss. In the agnostic setting, the maximum likelihood method fails as there may be no consistent hypothesis. We present a modified version of the maximum likelihood method, which still has the “consistency” component and adopts log output size as a regulation term. The algorithm operates as follows.
@@
    Finding a set of plausible hypotheses which make at most $m\cdot(r+\varepsilon)$ mistakes: For any hypothesis $g$, let $I_{g}=\{i|v_{i}\notin g(x_{i})\}$ denote the indices of training points that the hypothesis $g$ is inconsistent with. Then let $\widehat{\mathcal{H}}=\{g\in\mathcal{H}||I_{g}|\leq m\cdot(r+\varepsilon)\}$ be the set of hypotheses making at most $m\cdot(r+\varepsilon)$ mistakes.
@@
    Returning the hypothesis with low empirical log truncated output size: Return
@@
Given any $r\in[0,1]$ and $m\geq O(\frac{\log(|\mathcal{H}|)+\log(1/\delta)}{\varepsilon^{2}})$ IID training data, the modified maximum likelihood method can return a hypothesis satisfying
@@
Proof Proof Sketch Let $g^{\dagger}$ to be the hypothesis with recall loss at most $r$ and precision loss $p$. Let $\overline{r}=r+2\varepsilon$ and $\overline{p}=p+2\varepsilon$. Similar to the realizable setting, we build connection between precision and recall with empirical log output size (Lemma 2), i.e., for any constant $c\in(0,1]$,
$\widehat{\ell}^{\textrm{precision}}(g)\leq\frac{1+c}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+\frac{1+c}{c}\widehat{\ell}^{\textrm{recall}}(g)\,.$
@@
where $B=\{i|\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}\}$. The term (b) is lower bounded by the recall loss in Lemma 4. Intuitively, if $\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is small, the recall loss must be large while any hypothesis in $\widehat{\mathcal{H}}$ has a small empirical recall loss.
For term (a), since $\widehat{\ell}^{\textrm{recall}}(g^{\dagger})\leq\overline{r}$ and $\widehat{\ell}^{\textrm{precision}}(g^{\dagger})\leq\overline{p}$, at most training points, $\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is in $\frac{1}{2},2]$ (if it’s too large at $x_{i}$, precision loss is large; if it’s too small, recall loss is large). Also, for any hypothesis in $\widehat{\mathcal{H}}$, the empirical recall loss is small and thus, at most training points, $\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}$. At these training points satisfying $\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is in $\frac{1}{2},2]$ and $\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}$, we have
$\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$
$=$ $\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}+\log\frac{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$